OpenStudy (anonymous):

A boy weighing 150N runs a. up and b.)down a flight of stairs, the vertical distance between the floor is 5.5m . How much work is done by gravity?

OpenStudy (anonymous):

a-W= -mgh= (-150*5.5) since work is done against direction of gravity b- W= mgh =150*5.5 since work is done in direction of gravity

OpenStudy (anonymous):

how can i get tha angle of up and down?

OpenStudy (rane):

use cosine and sine ratio

OpenStudy (anonymous):

how i will compute the angle if the given was the force and h of the vertical distance of the floor?

OpenStudy (anonymous):

You cant. But you don't need to. Since gravity acts only in the vertical you don't need the angle or the horizontal distance to calculate your answer.

OpenStudy (anonymous):

thank you!